OpenStudy (anonymous):

Simplify this expression. √75+ √3 Can you please help me and explain how you got it? Medal for everyone that helps!!!

OpenStudy (henryblah):

Take 5√3. What do you get if you try and put the 5 back under the square root?

OpenStudy (anonymous):

15? I'm not really sure. I do homeschooling and my teacher doesn't help me.

OpenStudy (henryblah):

If you think about taking square no.s out of the square root. √75=√(25x3)=√25 x √3 = 5√3

OpenStudy (anonymous):

Is that the answer?

OpenStudy (anonymous):

not the final answer

OpenStudy (anonymous):

\[\sqrt{a} \times \sqrt{b} = \sqrt{ab}\] \[\sqrt{75} = \sqrt{25 \times 3} = \sqrt{25} \times \sqrt{3} = 5\sqrt{3}\] \[\sqrt{75} + \sqrt{3} = 5\sqrt{3} + \sqrt{3} = 6\sqrt{3}\]